# Limit theorems for a strongly irreducible product of independent random matrices under optimal moment assumptions

### v2

@@ 5.3. The bi-lateral pivoting technique
Let $(\gamma_{n})_{n\geq 0}$ be a random sequence in $\mathrm{End}(E)$. Let $(p_{k})_{k\geq 0}$ be a random sequence of integers such that $p_{2k+1}$ is equal to a non-random positive constant for all $k$. Let $(\tau_{k})_{k\geq 0}$ be independent an uniformly distributed and independent of the joint data of $\left(\widetilde{\gamma}^{p}_{k}\right)_{k\geq 0}$. Let $20<\varepsilon\leq 1/$ 2. Assume that points (2), (5) and (6) from Theorem 1.6 hold. For all $f\in E^{*}\cup\mathrm{End}(E)$, there exists a random integer $l^{f}\sim\mathcal{G}_{1/4}$, that is independent of $\left(\widetilde{\gamma}^{p}_{2k}\right)_{k\geq 0}$ and such that:
@@
Moreover, the sequence $(p_{2k+2}+m)_{k\geq 0}$ is i.i.d and $p_{2}$ has a finite exponential moment by Lemma 5.2, the sequence $\left(\overline{p}_{2k+2}-p_{0}-m\right)$ satisfies large deviations inequalities below the speed $-m-\mathbb{E}(p_{2})$ and by Lemma 5.3, the sequence $\left(\overline{p}_{2k+2}\right)$ also does. Therefore, there exits constants $C,\beta>0$ such that $\mathbb{P}\left(\overline{p}_{2k+2}\geq\alpha k\right)\leq Ce^{-\beta k}$ for all $k$. Let $C,\beta$ be such constants and $n\in\mathbb{N}$, for $k=\lfloor n/\alpha\rfloor$, we have:
@@
Now let $\alpha<\lambda=\lambda^{\prime}\lambda^{\prime\prime}$. Let $20<\delta<\lambda^{\prime\prime}/$ 2 and let $0<\alpha_{1}\leq\lambda^{\prime}$ be such that $\alpha<(\lambda^{\prime\prime}-2\delta)\alpha_{1}$. By the large deviations inequalities, there exist constants $C_{0},\beta_{0}$ such that the following assertions hold:
@@ 5.4. Quantitative estimates for the coefficients and spectral radius
Let us now give an upper bound on $-\log\left(\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}\right)$. By Definition of $l^{f}$ and $r_{n}^{v}$, we have $f\gamma^{p}_{0}\cdots\gamma^{p}_{2l^{f}}\mathbb{A}^{\varepsilon}\gamma_{2l^{f}+1}$ and $\gamma^{p}_{2q_{n}-2r_{n}-1}\mathbb{A}^{\varepsilon}\gamma_{\overline{p}_{2q_{n}-2r_{n}}}\cdots\gamma_{n-1}v$. Therefore, on the set $S_{n}^{f,v}$, by point (8) in Theorem 1.6, we have:
@@
(99) $-\log\left(\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}\right)\leq|\log(\varepsilon^{2}/4)|+\sum_{k=0}^{\overline{p}_{2l^{f}+1}-1}N(\gamma_{k})+\sum_{k=\overline{p}_{2q_{n}-2r_{n}^{v}}}^{n-1}N(\gamma_{k}).$
Let us write $I_{n}^{f,v}$ for the set $\{0,\dots,\overline{p}_{2l^{f}+1}-1\}\cup\{\overline{p}_{2q_{n}-2r_{n}^{v}},n-1\}$. Note that #$\#I_{n}^{f,v}=n-\overline{p}_{2q_{n}-2r_{n}^{v}}+\overline{p}_{2l^{f}+1}\geq 1$.
@@
$\leq\sum_{k=1}^{\infty}\mathbb{P}\left(\#I_{n}^{f,v}=k\right)2kN_{*}\nu\left(\frac{t-|\log(\varepsilon^{2}/4)|}{k}-B,+\infty\right)$
@@
Let us now give an upper bound on $-\log\left(\frac{\rho_{1}(\gamma_{n})}{\|\overline{\gamma}_{n}\|}\right)$. Let $S_{n}=\left(2c_{n}+1<q_{n}\right)$ By definition of $c_{n}$, and by point (8) in Theorem 1.6, on $S_{n}$ for:
@@
Outside of $S_{n}$, we have $-\log\left(\frac{\rho_{1}(\overline{\gamma}_{n})}{\|\overline{\gamma}_{n}\|}\right)\leq N(\overline{\gamma}_{n})\leq\sum_{k=0}^{n-1}N(\gamma_{k})$. Outside of $S_{n}$, let $J_{n}:=\{0,\dots,n-1\}$ and on $S_{n}$, let $J_{n}:=\{0,\dots,\overline{p}_{2c_{n}+1}-1\}\cup\{\overline{p}_{2q_{n}-2c_{n}-1},\dots,n-1\}$. Then, for all $t\geq 0$, we have:
@@ 5.5. Law of large numbers for the singular gap
This paragraph is dedicated to the proof of Theorem 1.7. Let $\nu$ be strongly irreducible and proximal on $\mathrm{End}(E)$. Let $\left(\gamma_{n}\right)_{n\geq 0}\sim\nu^{\otimes\mathbb{N}}$. We remind that Theorem 1.7 tells us that $\frac{-\log\sigma(\overline{\gamma}_{n})}{n}$ converges almost surely to a non-random positive limit that we denote by $\lambda_{1-2}(\nu)$ and (6) tells us that the random sequence $\left(-\log\sigma(\overline{\gamma}_{n})\right)_{n}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}(\nu)$. The proof of (6) is pretty straightforward from Lemma 5.8 and it implies that lim $\liminf\frac{-\log\sigma(\overline{\gamma}_{n})}{n}\geq\lambda_{1-2}(\nu)$ almost surely. The hard part is to prove that lim $\limsup\frac{-\log\sigma(\overline{\gamma}_{n})}{n}\leq\lambda_{1-2}(\nu)$. Indeed, we can not derive moment conditions on $N_{*}\nu$ from the fact that $\lambda_{1-2}(\nu)<+\infty$ only. For the proof of the almost sure convergence, we use the following Lemma.
@@
Let $\nu$ be a strongly irreducible and proximal probability distribution on $\mathrm{End}(E)$. Let $(\gamma_{n})\sim\nu^{\otimes\mathbb{N}}$. Let $20<\varepsilon\leq 1/$ 2 be as in Theorem 1.6. There exist constants $l_{0}\in\mathbb{N}$, and $10<\beta<$ 1 such that:
@@
Then we compose (108) with (107) and with the fact that $n_{1}$ exists with probability at least $\delta$ and we get (106). Now note that (106) combined with the fact that $\overline{p}_{2}<l_{1}$ outside of an event of probability $31-\beta^{2}\delta/$ 3 implies that:
@@
Moreover, we have $-\log\sigma(g_{1})>\lambda^{\prime\prime}n_{1}$ and $n_{1}\geq n_{0}$. Therefore:
@@ 5.6. Contraction property
Note also that $\Omega^{\prime}(E)\neq\Omega(E)$. For example let $E=\mathbb{R}^{2}$, and let $\pi_{1}$ and $\pi_{2}$ be the orthogonal projections onto the first and second coordinates. If $\gamma_{0}=\pi_{1}$ and $\gamma_{k}=\pi_{1}+2\pi_{2}$ for all $k\geq 1$, then $\overline{\gamma}_{n}=\pi_{1}$ for all $n\geq 1$ and $[\gamma_{k}\cdots\gamma_{n}]\underset{m\to\infty}{\longrightarrow}[\pi_{2}]$ for all $k\geq 1$. So $l^{\infty}(\gamma)$ is the first coordinate axis and $l^{\infty}(T\gamma)$ is the second coordinate axis. Hence $\gamma_{0}l^{\infty}(T\gamma)=[0]\neq l^{\infty}(\gamma)$ so $\gamma\in\Omega^{\prime}(E)\setminus\Omega(E)$.
@@
Let $u_{n}^{\varepsilon^{\prime}}\in U^{\varepsilon^{\prime}}\left(\overline{\gamma}_{n}\right)\setminus\{0\}$ for all $n\in\mathbb{N}$ and all $\varepsilon>0$. Note that (113) holds for all $0<\varepsilon^{\prime}<\varepsilon$ Therefore, for all $n_{0}\leq n\leq k$, we have $\mathrm{d}(u_{n},u_{k})\leq 4e^{-\alpha n}/\varepsilon^{\prime}$, by (113). Therefore $(u_{n}^{\varepsilon^{\prime}})_{n\geq 0}$ a Cauchy sequence on the full measure set $(n_{\alpha}<+\infty)$ so $(\gamma_{n})_{n\geq 0}\in\omega^{\prime}(E)$ on $(n_{\alpha}<+\infty)$ and therefore $\nu^{\otimes\mathbb{N}}(\Omega^{\prime}(E))=1$. Since (113) holds for all $\alpha<\lambda_{1-2}$, the random sequence $\left(-\log\mathrm{d}\left([u_{n}^{1}],l^{\infty}(\Omega)\right)\right)_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}$, which implies (109). When $v\notin\underline{\ker}(\nu)$, the random sequence $\left(-\log\mathrm{d}\left([\overline{\gamma}_{n}v],l^{\infty}\right)\right)_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}$, which implies (110). Then, we have $[\overline{\gamma}_{n}v]\to l^{\infty}(\gamma)$ and by the same reasoning applied to $(\overline{\gamma}_{n})_{n\geq 1}$, we have $[\gamma_{1}\cdots{\gamma}_{n}v]\to l^{\infty}(T\gamma)$. Hence, we have $\gamma_{0}l^{\infty}(T\gamma)=l^{\infty}$ almost surely so $\nu^{\otimes\mathbb{N}}(\Omega)=1$ By the same argument, for all $\alpha<\lambda_{1-2}$, there exist constants $C^{\prime},\beta^{\prime}>0$ such that:
@@
Now we prove the second part of (10), namely that the random sequence $(-\log\mathrm{d}(E^{+}(\overline{\gamma}_{n}),l^{\infty}(\gamma))_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}(\nu)$. Let $n\geq 0$ and assume that $2c_{n}+1<q_{n}$. By point 8, we have $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}=\gamma^{p}_{0}\cdots\gamma^{p}_{2q_{n}-2c_{n}-2}\mathbb{A}^{\frac{\varepsilon}{2}}f_{n}$. Then by Lemma 2.15, we have $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}\allowbreak(f_{n}z_{n})^{k}$ for all $k\geq 0$ moreover, for all choice of $e_{n}\in E^{+}(f_{n}z_{n})\setminus\{0\}$ and $e^{\prime}_{n}\in E^{+}(^{t}z_{n}{{}^{t}\!f_{n}})\setminus\{0\}$, we have $[(f_{n}z_{n})^{m}]\underset{k\to\infty}{\longrightarrow}[e_{n}e^{\prime}_{n}]$. Moreover $\mathbb{A}^{\varepsilon}$ is bi-homogeneous and closed so $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}e_{n}e^{\prime}_{n}$ and therefore $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}e_{n}$.

### v3

@@ 5.3. The bi-lateral pivoting technique
Let $(\gamma_{n})_{n\geq 0}$ be a random sequence in $\mathrm{End}(E)$. Let $(p_{k})_{k\geq 0}$ be a random sequence of integers such that $p_{2k+1}$ is equal to a non-random positive constant for all $k$. Let $(\tau_{k})_{k\geq 0}$ be independent an uniformly distributed and independent of the joint data of $\left(\widetilde{\gamma}^{p}_{k}\right)_{k\geq 0}$. Let $0<\varepsilon\leq 1/2$. Assume that points (2), (5) and (6) from Theorem 1.6 hold. For all $f\in E^{*}\cup\mathrm{End}(E)$, there exists a random integer $l^{f}\sim\mathcal{G}_{1/4}$, that is independent of $\left(\widetilde{\gamma}^{p}_{2k}\right)_{k\geq 0}$ and such that:
@@
Moreover, the sequence $(p_{2k+2}+m)_{k\geq 0}$ is i.i.d and $p_{2}$ has a finite exponential moment by Lemma 5.2, the sequence $\left(\overline{p}_{2k+2}-p_{0}-m\right)$ satisfies large deviations inequalities below the speed $m-\mathbb{E}(p_{2})$ and by Lemma 5.3, the sequence $\left(\overline{p}_{2k+2}\right)$ also does. Therefore, there exits constants $C,\beta>0$ such that $\mathbb{P}\left(\overline{p}_{2k+2}\geq\alpha k\right)\leq Ce^{-\beta k}$ for all $k$. Let $C,\beta$ be such constants and $n\in\mathbb{N}$, for $k=\lfloor n/\alpha\rfloor$, we have:
@@
Now let $\alpha<\lambda=\lambda^{\prime}\lambda^{\prime\prime}$. Let $0<\delta<\lambda^{\prime\prime}/2$ and let $0<\alpha_{1}\leq\lambda^{\prime}$ be such that $\alpha<(\lambda^{\prime\prime}-2\delta)\alpha_{1}$. By the large deviations inequalities, there exist constants $C_{0},\beta_{0}$ such that the following assertions hold:
@@ 5.4. Quantitative estimates for the coefficients and spectral radius
Let us now give an upper bound on $\log\left(\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}\right)$. By Definition of $l^{f}$ and $r_{n}^{v}$, we have $f\gamma^{p}_{0}\cdots\gamma^{p}_{2l^{f}}\mathbb{A}^{\varepsilon}\gamma_{2l^{f}+1}$ and $\gamma^{p}_{2q_{n}-2r_{n}-1}\mathbb{A}^{\varepsilon}\gamma_{\overline{p}_{2q_{n}-2r_{n}}}\cdots\gamma_{n-1}v$. Therefore, on the set $S_{n}^{f,v}$, by point (8) in Theorem 1.6, we have:
@@
(99) $\log\left(\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}\right)\leq|\log(\varepsilon^{2}/4)|+\sum_{k=0}^{\overline{p}_{2l^{f}+1}-1}N(\gamma_{k})+\sum_{k=\overline{p}_{2q_{n}-2r_{n}^{v}}}^{n-1}N(\gamma_{k}).$
Let us write $I_{n}^{f,v}$ for the set $\{0,\dots,\overline{p}_{2l^{f}+1}-1\}\cup\{\overline{p}_{2q_{n}-2r_{n}^{v}},n-1\}$. Note that $I_{n}^{f,v}=n-\overline{p}_{2q_{n}-2r_{n}^{v}}+\overline{p}_{2l^{f}+1}\geq 1$.
@@
$\leq\sum_{k=1}^{\infty}\mathbb{P}\left(\#I_{n}^{f,v}=k\right)2kN_{*}\nu\left(\frac{t-|\log(\varepsilon^{2}/4)|}{k}-B,+\infty\right)$ .
@@
Let us now give an upper bound on $\log\left(\frac{\rho_{1}(\gamma_{n})}{\|\overline{\gamma}_{n}\|}\right)$. Let $S_{n}=\left(2c_{n}+1<q_{n}\right)$ By definition of $c_{n}$, and by point (8) in Theorem 1.6, on $S_{n}$ for:
@@
Outside of $S_{n}$, we have $\log\left(\frac{\rho_{1}(\overline{\gamma}_{n})}{\|\overline{\gamma}_{n}\|}\right)\leq N(\overline{\gamma}_{n})\leq\sum_{k=0}^{n-1}N(\gamma_{k})$. Outside of $S_{n}$, let $J_{n}:=\{0,\dots,n-1\}$ and on $S_{n}$, let $J_{n}:=\{0,\dots,\overline{p}_{2c_{n}+1}-1\}\cup\{\overline{p}_{2q_{n}-2c_{n}-1},\dots,n-1\}$. Then, for all $t\geq 0$, we have:
@@ 5.5. Law of large numbers for the singular gap
This paragraph is dedicated to the proof of Theorem 1.7. Let $\nu$ be strongly irreducible and proximal on $\mathrm{End}(E)$. Let $\left(\gamma_{n}\right)_{n\geq 0}\sim\nu^{\otimes\mathbb{N}}$. We remind that Theorem 1.7 tells us that $\frac{-\log\sigma(\overline{\gamma}_{n})}{n}$ converges almost surely to a non-random positive limit that we denote by $\lambda_{1-2}(\nu)$ and (6) tells us that the random sequence $\left(-\log\sigma(\overline{\gamma}_{n})\right)_{n}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}(\nu)$. The proof of (6) is pretty straightforward from Lemma 5.8 and it implies that $\liminf\frac{-\log\sigma(\overline{\gamma}_{n})}{n}\geq\lambda_{1-2}(\nu)$ almost surely. The hard part is to prove that $\limsup\frac{-\log\sigma(\overline{\gamma}_{n})}{n}\leq\lambda_{1-2}(\nu)$. Indeed, we can not derive moment conditions on $N_{*}\nu$ from the fact that $\lambda_{1-2}(\nu)<+\infty$ only. For the proof of the almost sure convergence, we use the following Lemma.
@@
Let $\nu$ be a strongly irreducible and proximal probability distribution on $\mathrm{End}(E)$. Let $(\gamma_{n})\sim\nu^{\otimes\mathbb{N}}$. Let $0<\varepsilon\leq 1/2$ be as in Theorem 1.6. There exist constants $l_{0}\in\mathbb{N}$, and $0<\beta<1$ such that:
@@
Then we compose (108) with (107) and with the fact that $n_{1}$ exists with probability at least $\delta$ and we get (106). Now note that (106) combined with the fact that $\overline{p}_{2}<l_{1}$ outside of an event of probability $1-\beta^{2}\delta/3$ implies that:
@@
Moreover, we have $\log\sigma(g_{1})>\lambda^{\prime\prime}n_{1}$ and $n_{1}\geq n_{0}$. Therefore:
@@ 5.6. Contraction property
Note also that $\Omega^{\prime}(E)\neq\Omega(E)$. For example let $E=\mathbb{R}^{2}$, and let $\pi_{1}$ and $\pi_{2}$ be the orthogonal projections onto the first and second coordinates. If $\gamma_{0}=\pi_{1}$ and $\gamma_{k}=\pi_{1}+2\pi_{2}$ for all $k\geq 1$, then $\overline{\gamma}_{n}=\pi_{1}$ for all $n\geq 1$ and $\gamma_{k}\cdots\gamma_{n}]\underset{m\to\infty}{\longrightarrow}[\pi_{2}]$ for all $k\geq 1$. So $l^{\infty}(\gamma)$ is the first coordinate axis and $l^{\infty}(T\gamma)$ is the second coordinate axis. Hence $\gamma_{0}l^{\infty}(T\gamma)=[0]\neq l^{\infty}(\gamma)$ so $\gamma\in\Omega^{\prime}(E)\setminus\Omega(E)$.
@@
Let $u_{n}^{\varepsilon^{\prime}}\in U^{\varepsilon^{\prime}}\left(\overline{\gamma}_{n}\right)\setminus\{0\}$ for all $n\in\mathbb{N}$ and all $\varepsilon>0$. Note that (113) holds for all $0<\varepsilon^{\prime}<\varepsilon$ Therefore, for all $n_{0}\leq n\leq k$, we have $\mathrm{d}(u_{n},u_{k})\leq 4e^{-\alpha n}/\varepsilon^{\prime}$, by (113). Therefore $(u_{n}^{\varepsilon^{\prime}})_{n\geq 0}$ a Cauchy sequence on the full measure set $(n_{\alpha}<+\infty)$ so $(\gamma_{n})_{n\geq 0}\in\omega^{\prime}(E)$ on $(n_{\alpha}<+\infty)$ and therefore $\nu^{\otimes\mathbb{N}}(\Omega^{\prime}(E))=1$. Since (113) holds for all $\alpha<\lambda_{1-2}$, the random sequence $\left(-\log\mathrm{d}\left([u_{n}^{1}],l^{\infty}(\Omega)\right)\right)_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}$, which implies (109). When $v\notin\underline{\ker}(\nu)$, the random sequence $\left(-\log\mathrm{d}\left([\overline{\gamma}_{n}v],l^{\infty}\right)\right)_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}$, which implies (110). Then, we have $\overline{\gamma}_{n}v]\to l^{\infty}(\gamma)$ and by the same reasoning applied to $(\overline{\gamma}_{n})_{n\geq 1}$, we have $\gamma_{1}\cdots{\gamma}_{n}v]\to l^{\infty}(T\gamma)$. Hence, we have $\gamma_{0}l^{\infty}(T\gamma)=l^{\infty}$ almost surely so $\nu^{\otimes\mathbb{N}}(\Omega)=1$ By the same argument, for all $\alpha<\lambda_{1-2}$, there exist constants $C^{\prime},\beta^{\prime}>0$ such that:
@@
Now we prove the second part of (10), namely that the random sequence $(-\log\mathrm{d}(E^{+}(\overline{\gamma}_{n}),l^{\infty}(\gamma))_{n\geq 0}$ satisfies large deviations inequalities below the speed $\lambda_{1-2}(\nu)$. Let $n\geq 0$ and assume that $2c_{n}+1<q_{n}$. By point 8, we have $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}=\gamma^{p}_{0}\cdots\gamma^{p}_{2q_{n}-2c_{n}-2}\mathbb{A}^{\frac{\varepsilon}{2}}f_{n}$. Then by Lemma 2.15, we have $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}\allowbreak(f_{n}z_{n})^{k}$ for all $k\geq 0$ moreover, for all choice of $e_{n}\in E^{+}(f_{n}z_{n})\setminus\{0\}$ and $e^{\prime}_{n}\in E^{+}(^{t}z_{n}{{}^{t}\!f_{n}})\setminus\{0\}$, we have $(f_{n}z_{n})^{m}]\underset{k\to\infty}{\longrightarrow}[e_{n}e^{\prime}_{n}]$. Moreover $\mathbb{A}^{\varepsilon}$ is bi-homogeneous and closed so $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}e_{n}e^{\prime}_{n}$ and therefore $\overline{\gamma}^{p}_{2q_{n}-2c_{n}-1}\mathbb{A}^{\frac{\varepsilon}{4}}e_{n}$.